OpenStudy (anonymous):

find the equation of the tangent line y= √x/(x+1) (4,0.4)

OpenStudy (anonymous):

find the derivative replace \(x\) by 4 to get the slope use the point slope formula to find the equation for the line do you have the derivative?

OpenStudy (anonymous):

I keep getting confused when getting the derivative.

OpenStudy (anonymous):

quotient rule for this one \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\]

OpenStudy (anonymous):

no need to do much algebra to simplify, just make sure it is in the form where you can replace \(x\) by 4

OpenStudy (anonymous):

get \[\frac{(1+x)\frac{1}{2\sqrt{x}}-\sqrt{x}}{(x+1)^2}\] by direct substitution in to quotient rule

OpenStudy (anonymous):

i got -1/50, not sure if it's right

OpenStudy (anonymous):

let me see what i get

OpenStudy (anonymous):

i think i get something diffrerent

OpenStudy (anonymous):

i get \[\frac{\frac{5}{4}-2}{25}=-\frac{3}{100}\]

OpenStudy (anonymous):

check my arithmetic, but i think it is right

OpenStudy (anonymous):

yeah i got that the second time, but when i search it up on different websites ppl got like 1/50